you and your friend live 12 miles apart. You leave home at the same time and travel toward each other. You walk at a rate of 4 miles per hour and your friend bicycles at a rage of 11 mph. How long arter you leave home will you meet? How far from home will each of you be? i keep trying this problem and cannot come up with any answer - can you please help me? thank you leah
Answer by checkley71(8403)

12=4x+11x or 12=15x or x=12/15 or x=4/5 or x=.8 hours
.8*4=3.2 miles for the walker & .8*11=8.8 miles for the biker.
3.2+8.8=12 miles.